OpenStudy (anonymous):

Tell whether the line for each pair of equations are parallel, perpendicular, or neither. y = -3x + 7 -2x + 6y = 3

OpenStudy (anonymous):

Do you know how to make the second equation into y = mx + c?

OpenStudy (anonymous):

\[y = -3x + 7\] \[-2x + 6y = 3 \Rightarrow y = \frac{x}{3}+\frac{1}{2}\] Since we now have it in y = mx + c form graph it :).

OpenStudy (anonymous):

Okay thanks :D Lemme figure this out xD